Mr. Ayala borrows 100,000 at 10% effective annual interest. He must pay back the loan over 30 years with uniform monthly payments due on the 1st day of each month. What does Ayala pay each month? 


1
Expert's answer
2022-01-23T15:41:42-0500

Borrowed fund = 100,000

Effective annual interest = 10%

Loan period = 30 years

Monthly interest rate = 0.007974

Total number of payments = 360

Then, monthly payment (EMI) will be:


"EMI=100000\times \frac{0.007974 \times(1+0.007974)^{359}}{(1+0.007974)^{359}-1}"


"EMI=839.19"